OpenStudy (artfis):

Solve (f/g)(3) : Problem below HELP NEEDED!

OpenStudy (artfis):

\[2x^2-7\div8x-3\]

OpenStudy (shamim):

ur question is not clear to me

OpenStudy (artfis):

f(x) = \[2x^2 - 7\] g(x) = 8x - 3 f(x) divided over g(x) if possible, then multiplied altogether by 3. - DEVM105x Math

OpenStudy (artfis):

This is all combining functions type math. Operations with functions.

OpenStudy (shamim):

\[\frac{ f(x) }{ g(x) } \times 3 =\frac{ 3(2x ^{2-7}) }{ 8x-3 }\]

OpenStudy (artfis):

\[\frac{ 6x^2-21 }{ 8x-3 } ?\]

OpenStudy (shamim):

ya

OpenStudy (artfis):

Gotcha! Thank you so much